On analysis, a coal sample has the following composition by weight; C = 90%, O = 4%, S


= 0.5% and ash = 5%. Net calorific value was found to be 8,480 kcal/kg. Calculate the


percentage of hydrogen and gross calorific value of coal.

1
Expert's answer
2022-03-07T17:17:06-0500

C = 90%

O = 4%

S = 0.5%

ash = 5%

Net calorific value (LCV) = 8480 kcal/kg


Solution:

Dulong’s formula for calculating the calorific value can be used:

Gross calorific value (HCV) = (1/100)×[8080C + 34500(H − O/8) + 2240S],

where C, S, O, H are the percentages of carbon, sulphur, oxygen and hydrogen.

Net calorific value (LCV) = HCV − 0.09H×587


LCV = HCV − 0.09H×587

LCV = (1/100)[8080C + 34500(H − O/8) + 2240S] − 0.09H×587

8480 = (0.01) × [8080×90 + 34500×(H − 4/8) + 2240×0.5] − 0.09H×587

8480 = 7272 + 345H − 172.5 + 11.2 − 52.83H

1369.3 = 292.17H

H = (1369.3) / (292.17) = 4.687%

H = 4.687%


HCV = (1/100)×[8080C + 34500(H − O/8) + 2240S]

HCV = (0.01) × [8080×90 + 34500×(4.687 − 4/8) + 2240×0.5] = 8727.715

Gross calorific value (HCV) = 8727.715 kcal/kg


Answer:

The percentage of hydrogen (H) is 4.687%

The gross calorific value of coal (HCV) is 8727.715 kcal/kg